Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 05

Figure 1: Brief answers to selected problems in section 2.6, part 1 Figure 2: Brief answers to selected problems in section 2.6, part 2

2 Figure 3: Brief answers to selected problems in section 3.2

3 Figure 4: Brief answers to selected problems in section 3.3

4